# Wigner’s friend scenarios: on what to condition and how to verify the predictions

### v1

@@ VII Conclusions
Finally, we comment on the memory loss of the aforementioned friend whose laboratory, including his memory, underwent the measurement procedure. It is sometimes argued that because of the loss of memory, the person should not be considered the same throughout the experiment (see [14] and references therein). In such a case, the individual would not be considered the same “friend” or “player” in our game. Consequently, they might not accept that they made a prediction in the past that led to them losing the game. The conventional notion of personal identity is defined by the distinctive ways in which an individual defines themselves, encompassing aspects such as self-image, cultural references, political credos, or other interests. Since the loss of memory in the measurement solely affects the recollection of the outcome of the very last measurement, these fundamental aspects of identity are expected to remain unaltered. Moreover, it is not uncommon for individuals to forget events from the past. However, when presented with compelling evidence that an event occurred in the past, they tend to accept that it did occur and that they simply forgot about it. Similarly, in our game, after players are presented with the signed paper containing their prediction, they may accept it as legitimate. Finally, one might approach the problem of identity in the present case pragmatically: Would you enter the game under the risk of losing the memory of one bit of information, if the award were one million coins?
@@ Acknowledgements
FDS and ČB acknowledge the discussion with Veronika Baumann. This research was funded in whole or in part by the Austrian Science Fund (FWF) 10.55776/COE1 (Quantum Science Austria), [10.55776/F71] (BeyondC) and [10.55776/RG3] (Reseacrh Group 3). For Open Access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. This publication was made possible through the financial support of the ID 62312 grant from the John Templeton Foundation, as part of The Quantum Information Structure of Spacetime (QISS) Project (qiss.fr). The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. We also acknowledge support from CoQuSy project PID2022-140506NB-C21, and the María de Maeztu project CEX2021-001164-M for Units of Excellence, funded by MICIU/AEI/10.13039/501100011033/FEDER, UE. GM acknowledges the Ramón y Cajal program RYC2021-031121-I funded by MICIU/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR. This research was also supported by the FWF (Austrian Science Fund) through an Erwin Schrödinger Fellowship (Project J 4699)
@@ Appendix A General encoding for the transmitted information between labs
In this Appendix we consider generic scenarios allowing that information about the qubit measurement results in the friend’s laboratory becomes accessible to Wigner. More precisely, we consider unitary interactions that partially correlate the accessible environmental states in the friend’s laboratory $\{|\phi_{0}\rangle_{F},|\phi_{1}\rangle_{F}\}$ (containing univocal information about the outcome of the qubit measurement) with those in Wigner’s laboratory $\{|\omega_{0}\rangle_{W},|\omega_{1}\rangle_{W}\}$. This can be achieved by generalizing the unitary that couples the system, the friend’s lab and Wigner’s environment defined through Eq. (4) by allowing it to depend on a parameter $\theta$, i.e. $U_{\mathrm{int}}(\theta)$ with $\theta\in[0,1]$, so that its relevant action over the states of the system, the friend and Wigner’s environments is now:
@@
Here we introduced the following parameterized states of the system and the friend’s lab:
@@
which reduce to the two standard Bell states given below Eq. (4) for $\theta=0$. Henceforth it is easy to check that the above unitary $U_{\mathrm{int}}(\theta)$ reduces to Eq. (4) for the case $\theta=0$. On the other hand, for $\theta>0$ such a unitary generically encodes information about the qubit measurement results in friend’s lab into superpositions of the visible states of Wigner’s environment with different amplitudes that depend on $\theta$. In the extreme case $\theta=1$ we have $|\Phi_{1}^{+}\rangle_{W}=|\uparrow_{0}\rangle_{S}|\phi_{0}\rangle_{F}$ and $|\varphi_{1}^{-}\rangle_{W}=|\downarrow_{0}\rangle_{S}|\phi_{1}\rangle_{F}$ (while $|\Phi_{1}^{-}\rangle=|\varphi_{1}^{+}\rangle=0$) so that Wigner’s accessible lab states become perfectly correlated with the measured qubit states and friend’s lab results.
Applying $U_{\mathrm{int}}(\theta)$ to Wigner’s state in Eq. (2) we obtain [c.f. Eq. (5)] :
@@
Therefore, depending on the value of $\theta$, Wigner’s environmental states $|\omega_{0}\rangle_{E}$ and $|\omega_{1}\rangle_{E}$ can respectively get correlated to the system the and friend’s lab states $|\uparrow\rangle_{S}|\phi_{0}\rangle_{F}$ and $|\downarrow\rangle_{S}|\phi_{1}\rangle_{F}$. In particular, the correlation is stronger, the larger the value of $\theta$ is, while for $\theta=0$ the two last lines of the state above disappear and we recover Eq. (5).
@@
with perfect correlations between the preferred basis states of the friend’s and Wigner lab degrees of freedom. Therefore the information about the friend’s measurement outcomes is completely accessible from Wigner’s environmental states without ambiguity in this case.
@@ Appendix B Wigner’s and friend’s predictions for partial leakage of information
Let’s now discuss the results of the checking procedure presented in Secs. IV and V for measurements in Wigner’s and friend’s bubbles, respectively, in the case of general encoding of information presented in Appendix A.
@@
with respective probabilities $p_{\omega_{0}}^{W}=[1+\cos(\pi\theta/2)]/2$ and $p_{\omega_{1}}^{W}=[1-\cos(\pi\theta/2)]/2$, which are submitted to the referee $R_{W}$. Notice that now the prediction of Wigner is not deterministic, but biased in general towards $\omega_{0}$, as $1/2\leq p_{\omega_{0}}^{W}\leq 1$ and $20\leq p_{\omega_{1}}^{W}\leq 1/$ 2. On the other hand, the friend’s prediction of the probabilities for observing one or the other outcome of MW, basing on state assignment Eq. (19), are again $p^{F}_{\omega_{0}}=1/2$ and $p^{F}_{\omega_{1}}=1/2$, which are also submitted to $R_{W}$. Notice that if no information about which outcome is present in the encoding, i.e. $\theta=0$, this reduces to the case discussed in Sec. IV. However, we observe that even if a part of the “which-outcome” information is leaked from the friend’s to Wigner’s environment, their predictions of the estimated probabilities of the two observers will still differ and hence the referee $R_{W}$ by performing the measurement $M_{W}$ can detect this mismatch performing a sufficient number of runs. Interestingly Wigner’s and the friend’s predictions become closer to each other as more information about the qubit measurement is leaked in the interaction between labs (which corresponds to increasing values of $\theta$) reaching the case of equal predictions for $\theta=1$.
@@
In particular, assuming Wald’s sequential probability ratio test [31], the referee $R_{W}$ may collect statistics from consecutive runs of the game and construct the cumulative log-likelihood ratio between Wigner’s and friend predictions. After $n$ runs obtaining outcomes $\{r_{1},r_{2},...,r_{n}\}$ the cumulative log-likelihood ratio would read:
@@
A similar situation is obtained in the case of performing the measurement $M_{F}$ in the friend’s bubble. Similarly to the situation described in Sec. V when a second measurement of the qubit state is performed in the friend’s lab the results will confirm the initial measurement state assignment: she predicts $p^{F}_{\phi_{0}}=1$ (or 0), $p^{F}_{\phi_{1}}=0$ (or 1) and gives this prediction to referee RF. On the other hand, from the point of view of Wigner, the state after the repeated measurement using unitary $U_{\rm meas}$ between system and an extra degree of freedom of the friend’s lab $F^{\prime}$ reads:

### v2

@@ VI Conclusions
Finally, we comment on the memory loss of the aforementioned friend whose laboratory, including his memory, underwent the measurement procedure. It is sometimes argued that, because of the loss of memory, the person should not be considered the same throughout the experiment (see [14] and references therein for a discussion). In such a case, the individual would not be considered the same “friend” or “player” in our game. Consequently, they might not accept that they made a prediction in the past that led to them losing the game. The conventional notion of personal identity is defined by the distinctive ways in which an individual defines themselves, encompassing aspects such as self-image, cultural references, political credos, or other interests. Since the loss of memory in the measurement solely affects the recollection of the outcome of the very last measurement, these fundamental aspects of identity are expected to remain unaltered. Moreover, it is not uncommon for individuals to forget events from the past. However, when presented with compelling evidence that an event occurred in the past, they tend to accept that it did occur and that they simply forgot about it. Similarly, in our game, after players are presented with the signed paper containing their prediction, they may accept it as legitimate. In the end, one might approach the problem of identity in the present case pragmatically: Would you enter the game under the risk of losing the memory of one bit of information, if the award were one million coins?
@@ Acknowledgements
FDS and ČB acknowledge the discussion with Veronika Baumann. This research was funded in whole or in part by the Austrian Science Fund (FWF) 10.55776/COE1 (Quantum Science Austria), [10.55776/F71] (BeyondC) and [10.55776/RG3] (Reseacrh Group 3). For Open Access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission. This publication was made possible through the financial support of the ID 62312 grant from the John Templeton Foundation, as part of The Quantum Information Structure of Spacetime (QISS) Project (qiss.fr). The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the John Templeton Foundation. We also acknowledge support from CoQuSy project PID2022-140506NB-C21, and the María de Maeztu project CEX2021-001164-M for Units of Excellence, funded by MICIU/AEI/10.13039/501100011033/FEDER, UE. GM acknowledges the Ramón y Cajal program RYC2021-031121-I funded by MICIU/AEI/10.13039/501100011033 and European Union NextGenerationEU/PRTR. This research was also supported by the FWF (Austrian Science Fund) through an Erwin Schrödinger Fellowship (Project J 4699).
@@ Appendix A General encoding for the transmitted information between labs
In this Appendix we consider generic scenarios allowing that some information about the qubit measurement results in the friend’s laboratory becomes accessible to Wigner. More precisely, we consider unitary interactions that partially correlate the accessible environmental states in the friend’s laboratory $\{|\phi_{0}\rangle_{F},|\phi_{1}\rangle_{F}\}$ (containing univocal information about the outcome of the qubit measurement) with those in Wigner’s laboratory $\{|\omega_{0}\rangle_{W},|\omega_{1}\rangle_{W}\}$. This can be achieved by generalizing the unitary that couples the system, the friend’s lab and Wigner’s environment defined through Eq. (4) by allowing it to depend on a parameter $\theta$, i.e. $U_{\mathrm{int}}(\theta)$ with $\theta\in[0,1]$, so that its relevant action over the states of the system, the friend and Wigner’s environments is now:
@@
Here we introduced the following (unnormalized) parameterized states of the system and the friend’s lab:
@@
with $\langle\Phi^{\pm}|\Phi^{\pm}\rangle=\langle\varphi^{\mp}|\varphi^{\mp}\rangle=[1\pm\sin(\pi\theta/2)]$, which reduce to the two standard Bell states given below Eq. (4) for $\theta=0$. Henceforth it is easy to check that the above unitary $U_{\mathrm{int}}(\theta)$ reduces to Eq. (4) for the case $\theta=0$. On the other hand, for $\theta>0$ such a unitary generically encodes partial information about the qubit measurement results in friend’s lab into superpositions of the states of Wigner’s environment with different amplitudes that depend on $\theta$. In the extreme case $\theta=1$ we have $|\Phi_{1}^{+}\rangle_{W}=|\uparrow_{0}\rangle_{S}|\phi_{0}\rangle_{F}$ and $|\varphi_{1}^{-}\rangle_{W}=|\downarrow_{0}\rangle_{S}|\phi_{1}\rangle_{F}$ (while $|\Phi_{1}^{-}\rangle=|\varphi_{1}^{+}\rangle=0$) so that Wigner’s accessible lab states become perfectly correlated with the measured qubit states and friend’s lab results.
Applying $U_{\mathrm{int}}(\theta)$ to Wigner’s state in Eq. (2) we obtain the unnormalized (c.f. Eq. (5)):
@@
Therefore, depending on the value of $\theta$, Wigner’s environmental states $|\omega_{0}\rangle_{E}$ and $|\omega_{1}\rangle_{E}$ can respectively get correlated to the system and the friend’s lab states $|\uparrow\rangle_{S}|\phi_{0}\rangle_{F}$ and $|\downarrow\rangle_{S}|\phi_{1}\rangle_{F}$. In particular, the correlation is stronger, the larger the value of $\theta$ is, while for $\theta=0$ the two last lines of the state above disappear and we recover Eq. (5).
@@
with perfect correlations between the preferred basis states of the friend’s and Wigner lab degrees of freedom. Therefore the information about the friend’s measurement outcomes is completely accessible from Wigner’s environmental states.
@@ Appendix B Wigner’s and friend’s predictions for partial leakage of information
Let’s now discuss the results of the checking procedure presented in Secs. IV.2 and IV.3 for measurements in Wigner’s and friend’s bubbles, respectively, in the case of general encoding of information presented in Appendix A.
@@
with respective probabilities $p_{\omega_{0}}^{W}=[1+\cos(\pi\theta/2)]/2$ and $p_{\omega_{1}}^{W}=[1-\cos(\pi\theta/2)]/2$, which are submitted to the referee $R_{W}$. Notice that now the prediction of Wigner is not deterministic, but biased in general towards $\omega_{0}$, as $1/2\leq p_{\omega_{0}}^{W}\leq 1$ and $20\leq p_{\omega_{1}}^{W}\leq 1/$ 2. On the other hand, the friend’s prediction of the probabilities for observing one or the other outcome of $M_{W}$, basing on state assignment Eq. (19), are again $p^{F}_{\omega_{0}}=1/2$ and $p^{F}_{\omega_{1}}=1/2$, which are also submitted to $R_{W}$. Notice that if no information about which outcome is present in the encoding, i.e. $\theta=0$, this reduces to the case discussed in Sec. IV.2. However, we observe that even if a part of the “which-outcome” information is leaked from the friend’s to Wigner’s environment, their predictions of the estimated probabilities of the two observers will still differ and hence the referee $R_{W}$ by performing the measurement $M_{W}$ can detect this mismatch performing a sufficient number of runs. Interestingly Wigner’s and the friend’s predictions become closer to each other as more information about the qubit measurement is leaked in the interaction between labs (which corresponds to increasing values of $\theta$) reaching the case of equal predictions for $\theta=1$.
@@
In particular, we consider Wald’s sequential probability ratio test [34], one of the three standard approaches to hypothesis testing. There, two hypotheses for the probability of a sequence of observed outcomes $r$ (here $p_{r}^{W}$ and $p_{r}^{F}$ with $r=\{\omega_{0},\omega_{1}\}$) are compared by constructing their log-likelihood ratio and computing its evolution as the number of observations increases. One of the two hypotheses is accepted if the log-likelihood reaches predefined positive or negative thresholds. In our situation, the referee $R_{W}$ may collect statistics from consecutive runs of the game and construct the cumulative log-likelihood ratio between Wigner’s and friend predictions. After $n$ runs obtaining outcomes $\{r_{1},r_{2},...,r_{n}\}$ the cumulative log-likelihood ratio would read:
@@
A similar situation is obtained in the case of performing the measurement $M_{F}$ in the friend’s bubble. Similarly to the situation described in Sec. IV.3 when a second measurement of the qubit state is performed in the friend’s lab the results will confirm the initial measurement state assignment: she predicts $p^{F}_{\phi_{0}}=1$ (or 0), $p^{F}_{\phi_{1}}=0$ (or 1) and gives this prediction to referee RF. On the other hand, from the point of view of Wigner, the state after the repeated measurement using unitary $U_{\rm meas}$ between system and an extra degree of freedom of the friend’s lab $F^{\prime}$ reads: